OpenStudy (anonymous):

Calculate the speed of an electron that has a relativistic factor γ = 1.2; express your answer in m/s as well as relative to c.

OpenStudy (noelgreco):

\[m _{r}=\frac{ m }{ \sqrt{1-\frac{ v ^{2} }{ c ^{2} }} }\]